OpenStudy (kobeni-chan):

MeDaL! :D Phillip received 75 points on a project for school. He can make changes and receive two-tenths of the missing points back. He can do this 10 times. Create the formula for the sum of this geometric series, and explain your steps in solving for the maximum grade Phillip can receive.

OpenStudy (littlebird):

The geometric sequence of the missing points is: a is the missing points initially. Assuming 100 marks, this is 25 r = 8/10, as the missing marks reduces by 2/10 each time u_n = 25( \frac{8}{10})^{n-1} Assuming the marks are out of 100, and marks are rounded to the nearest whole number, the maximum mark he can get is 98, as 2/10 of the remaining 2 marks would make it 98.4, which would round back to 98. This series is converging, as the 8/10 gets smaller every time, so the series will eventually converge.

OpenStudy (kobeni-chan):

Ok I think I get it. Thank you :)